OpenStudy (anonymous):

Medal to helper! Can you please check my answers? Important! 5Q's. Thanks!

OpenStudy (anonymous):

1. What is the slope of the line which passes through (−2, 0) and (0, 4)? A. −2 B. 0 C. 2 <--- My Answer D. Undefined

OpenStudy (anonymous):

2. Use the table below to answer this question: x y −1 7 3 3 5 1 Find the average rate of change for the given function from x = −1 to x = 5. A. −6 B. −1 <--- My Answer C. 1 D. 6

OpenStudy (anonymous):

3. What is the equation in point−slope form of the line passing through (0, 6) and (1, 3)? A. (y − 3) = −3(x − 1) <--- My Answer B. (y + 3) = 3(x + 1) C. (y + 3) = −3(x + 1) D. (y − 3) = 3(x + 6)

OpenStudy (anonymous):

4. What is the equation of the graph below? A. y = − (x + 2)^2 + 2 B. y = − (x − 3)^2 + 2 C. y = (x − 2)^2 + 2 D. y = (x + 3)^2 + 2 <--- My Answer

OpenStudy (anonymous):

5. Simplify (x − 4)(x^2 + 3x + 2). A. x^3 − x^2 + 14x − 8 B. x^3 + 7x^2 − 10x − 8 C. x^3 − x^2 − 10x − 8 D. x^3 + 7x^2 + 14x − 8 <--- My Answer

OpenStudy (anonymous):

@phi @Hero @SolomonZelman @tom982

OpenStudy (solomonzelman):

#1 Correct #2 Correct #3 Wrong #4 Wrong #5 Wrong

OpenStudy (solomonzelman):

No, #3 might be right, let me check again

OpenStudy (solomonzelman):

(0, 6) and (1, 3) slope=m=(6-3)/(0-1)=3/(-1)=-3 So, with slope -3 and point (1,3) you have: y - 3 = -3•(x-1) So #3 is correct. Alternatively, you could have used the other point (0,6) y - 6 = -3•(x-0) y - 6 = -3x

OpenStudy (solomonzelman):

yes, #3 is correct, but #4 and #5 are for sure wrong.

OpenStudy (anonymous):

Okay thanks. Could you help me with those two?

OpenStudy (solomonzelman):

Yes, surely. the general form of the equation of the parabola with vertex at (h,k) is: \(\large\color{#000000 }{ \displaystyle y=a(x-h)^2+k }\) If the parabola opens up, a>0 if the parabola opens down, a<0

OpenStudy (solomonzelman):

So, your coefficient must be negative, because the parabola opens down.

OpenStudy (solomonzelman):

And, tell me where is the center of your parabola located?

OpenStudy (solomonzelman):

I mean the vertex, not the center

OpenStudy (anonymous):

(3,2)

OpenStudy (solomonzelman):

Yes, that is right.

OpenStudy (anonymous):

So, y = − (x − 3)^2 + 2 is correct, right?

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y=a(x-h)^2+k }\) h=3 k=2 \(\large\color{#000000 }{ \displaystyle y=a(x-3)^2+2 }\) So the closes option is \(\large\color{#000000 }{ \displaystyle y=-(x-3)^2+2 }\)

OpenStudy (anonymous):

Alright thank you. Now number 5.

OpenStudy (solomonzelman):

If you want, also, you can solve for "a", and get the answer without eliminating the rest of the options. Want to solve for a, or to proceed to #5 right away?

OpenStudy (anonymous):

I would like to move on to #5.

OpenStudy (solomonzelman):

Alright ...

OpenStudy (solomonzelman):

You have to expand: \(\large\color{#000000 }{ \displaystyle (x-4)(x^2+3x+2) }\)

OpenStudy (solomonzelman):

Can you tell me what do you get if you were to expand \(\large\color{#000000 }{ \displaystyle (x-4)\cdot x^2 }\)

OpenStudy (solomonzelman):

You can use the following rules \(\large\color{#000000 }{ \displaystyle (a+b)\times c=(a\times c)+ (b\times c) }\) \(\large\color{#000000 }{ \displaystyle (a-b)\times c=(a\times c)- (b\times c) }\)

OpenStudy (anonymous):

multiply x^2 and x, then multiply x^2 and -4?

OpenStudy (solomonzelman):

yes, and wha do you get after doing this?

OpenStudy (anonymous):

x^3 - 4x^2

OpenStudy (solomonzelman):

yes, very good

OpenStudy (solomonzelman):

Now, what do you get after expanding \(\large\color{#000000 }{ \displaystyle (x-4)\times 3x }\)

OpenStudy (anonymous):

3x^2 - 12x

OpenStudy (solomonzelman):

Yes, very good

OpenStudy (solomonzelman):

And lastly what do you get after expanding \(\large\color{#000000 }{ \displaystyle (x-4)\times 2 }\)

OpenStudy (anonymous):

2x - 8

OpenStudy (solomonzelman):

Yup, now we will put the peaces together...

OpenStudy (solomonzelman):

We wanted to expand \(\large\color{#000000 }{ \displaystyle (x-4)(x^2+3x+2) }\) And this is the same as expanding \(\large\color{#000000 }{ \displaystyle (x-4)\cdot x^2 }\) \(\large\color{#000000 }{ \displaystyle (x-4)\cdot3x }\) \(\large\color{#000000 }{ \displaystyle (x-4)\cdot2 }\) separately, and then adding the results. We get that: \(\large\color{#000000 }{ \displaystyle (x-4)\cdot x^2=\quad +x^3-4x^2 }\) \(\large\color{#000000 }{ \displaystyle (x-4)\cdot3x=\quad +3x^2-12x }\) \(\large\color{#000000 }{ \displaystyle (x-4)\cdot2 =~{{\tiny~~}}\quad +2x-8}\) And add the results: \(\large\color{#000000 }{ \displaystyle x^3-4x^2+3x^2-12x+2x-8 }\)

OpenStudy (solomonzelman):

Can you simplify \(\large\color{#000000 }{ \displaystyle x^3-4x^2+3x^2-12x+2x-8 }\) ?

OpenStudy (anonymous):

Yes, let me try.

OpenStudy (solomonzelman):

go ahead and take your time ;)

OpenStudy (anonymous):

x^3 - x^2 - 10x - 8

OpenStudy (solomonzelman):

yup, and that is your final answer.

OpenStudy (anonymous):

Wow. I would have never guessed that one. You're explanation makes sense. Thank you so much!

OpenStudy (anonymous):

Could you check this too?: Which of the following is the conjugate of a complex number with 2 as the real part and −8 as the imaginary part? A. −2 + 8i B. 2 + 8i <--- My Answer C. 2 − 8i D. −2 − 8i

OpenStudy (solomonzelman):